OpenStudy (anonymous):

Integration int x/(2x+1)dx

OpenStudy (raden):

u = 2x + 1 ----> x = (u-1)/2 du = 2 dx dx = 1/2 du see the integral becomes : 1/2 int (u-1)/(2u) du = 1/4 int (u-1)/u du = 1/4 int (1 - 1/u) du conitinue it .....

OpenStudy (raden):

that should be easier now

OpenStudy (anonymous):

Thank you

OpenStudy (raden):

wlcm